### v1

Abstract
We consider routing in reconfigurable networks, which is also known as coflow scheduling in the literature. The algorithmic literature generally (perhaps implicitly) assumes that the amount of data to be transferred is large. Thus the standard way to model a collection of requested data transfers is by an integer demand matrix $D$, where the entry in row $i$ and column $j$ of $D$ is an integer representing the amount of information that the application wants to send from machine/node $i$ to machine/node $j$. A feasible coflow schedule is then a sequence of matchings, which represent the sequence of data transfers that covers $D$. In this work, we investigate coflow scheduling when the size of some of the requested data transfers may be small relative to the amount of data that can be transferred in one round. fractional matchings and/or that employ indirect routing, and compare the relative utility of these options. We design algorithms that perform much better for small demands than the algorithms in the literature that were designed for large data transfers.
1 Introduction
We consider routing in reconfigurable networks—also known as coflow scheduling—whose topologies dynamically adapt to the structure of the communication traffic [CS12a, AS25a]. There is currently no consensus in the community on what the best reconfigurable network design is, and the community lacks models and metrics to rigorously study and compare different designs [AS25a]. Our high level goal in this work is to fill some of this gap.
1.1 Background and Motivation
In coflow scheduling, there is a layer, which we will call the coflow layer, that sits underneath the application layer, and above the network layer, in the network protocol stack. The application passes the coflow protocol a collection of requests for data transfers that the application wants to execute. The task of the coflow layer is to sensibly schedule/manage when these requests for data transfers are passed to the networking layer. Most notably the coflow layer should not overwhelm the network with more requests than the network can simultaneously support.
The algorithmic literature in this area generally models this informal goal by the formal constraint that at all times the pairs of nodes communicating form a matching $M$, which in this context is a collection $M=\{(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{n},t_{n})\}$ of source-destination pairs of nodes with distinct sources and distinct sinks (so $i\neq j$ implies $s_{i}\neq s_{j}$ and $t_{i}\neq t_{j}$). A coflow schedule is then a sequence of matchings $M_{1},\ldots,M_{T}$, where matching $M_{t}$ is passed to the networking layer at time $t$, and during the time interval $[t,t+1]$ the network layer transmit a unit of data from $s_{i}$ to $t_{i}$ for each $(s_{i},t_{i})\in M_{t}$.
In the algorithmic literature on coflow scheduling [CS15a, CZS14a, Ahm+20a, Cho+19a, Im+19a, RS25a, JKR17a, QSZ15a, DM20a] the input passed to the coflow layer is assumed to consist of an integer demand matrix $D$, where $D_{ij}$, the entry in row $i$ and column $j$ of $D$, is an integer representing the amount of information that the application wants to send from machine/node $i$ to machine/node $j$. Then generally the objective is to optimize some standard quality-of-service metric, like minimizing makespan, which is the time when the last unit of data is transmitted, or minimizing the average completion (or delay) time, which is the average time that it takes a unit of data to reach its destination. So, in some sense, this is modeling scheduling coflows for a single application. (Some of the coflow scheduling literature also considers generalizations of this basic setting, see Section˜1.3.).
The algorithmic literature generally (perhaps implicitly) assumes that the number of requests for data transfers, and the size of these transfers, is large, as this was the situation that was the original motivation for the proposal of a coflow layer [CS12a]. This is reflected in the modeling assumption that the demands $D_{ij}$ are integer. The goal of our research was to examine coflow scheduling when one removes the assumption in the algorithmic coflow literature that the amount of data requested to be transferred in each requested data transfer is large. This assumption is not appropriate in settings where many of the data transfers are small, and it is certainly at least conceivable that such a situation could sometimes arise in data-parallel applications. We model this by allowing the demands $D_{ij}$ to be arbitrary nonnegative rational numbers, and in particular a demand $D_{ij}$ may be less than 1.
To understand the additional issues that arise if some of the requested data transfers are small, consider the demand matrix $D$ where each entry $D_{ij}=\frac{1}{2n^{3/2}}$, where $n$ is the number of nodes used by the application. Then these data transfers can be supported by routing up to a unit of data from node $i$ to node $j$ at the time $t$ where $t=(j-i)\mod n$. This results in a makespan of n1n-1 (and average completion/delay time of (n1)/2(n-1)/2). And it might seem that obtaining sublinear makespan is impossible, as $D$ cannot be expressed as a linear combination of less than n1n-1 matchings.
However, makespan $O(\sqrt{n})$ can be achieved if the coflow manager can indirectly route data (which is commonly assumed to be allowable in the reconfigurable network routing literature [Shr+19a, Bal+20a, Ami+24a, Gho+17a, Mel+24a, AAS23a]. As an example of indirect routing, a unit of data with source $s_{i}$ and destination $t_{i}$ could be routed using two matchings $M_{1}$ and $M_{2}$. So $M_{1}$ could send this data from source node $s_{i}$ to an intermediate node $k$, and $M_{2}$ could route this data from node $k$ to destination node $t_{i}$. One can achieve $O(\sqrt{n})$ makespan (for the instance where each $D_{ij}=\frac{1}{2n^{3/2}}$) in the following way. The $n$ nodes are conceptually assigned to nodes in an $\sqrt{n}$ by $\sqrt{n}$ grid square $G$. Then the data transfers are divided into two phases, and each phase is divided into $\sqrt{n}-1$ subphases. In subphase $k$ of Phase 1, the node in row $i$ column $j$ of $G$ routes to the node in row $(i+k)\mod\sqrt{n}$ column $j$ all its data whose final destination is in row $(i+k)\mod\sqrt{n}$. Thus after the first phase, the current location of each unit of data is in the same row as its final destination. Then subphase $k$ of Phase 2, the node in row $i$ column $j$ of $G$ routes to the node in row $i$ column $(j+k)\mod\sqrt{n}$ all its data whose final destination is that node. See Figure˜1 for an illustration of this algorithm when $n=9$.
Figure 1: The data transmissions for nodes in positions $(0,0)$ and $(1,1)$ for the 3 by 3 grid formed from $n=9$ nodes. The matchings $M_{1}$ and $M_{2}$ belong to Phase 1, and the matchings $M_{3}$ and $M_{4}$ belong to Phase 2.
Here we consider eight variations of coflow scheduling with rational demand matrices, where eight arises from there being two options to each of the following three characteristics:
Direct vs. Indirect Routing.
We consider both allowing and disallowing indirect transfers.
Makespan vs. Average Completion Times Objective.
We consider both of the two most commonly considered quality-of-service objectives, makespan and average completion time (which are essentially the infinity norm and one norm of the delays for the data units).
Fractional vs. Integral Routing.
One conceptual “solution” for coflow scheduling of small data transfers is to just change the assumptions about the networking layer to assuming that the network layer can route a fractional matching in each unit of time. In this context, a fractional matching is given by a collection $\{(s_{1},t_{1},p_{1}),(s_{2},t_{2},p_{2}),\ldots,(s_{n},t_{n},p_{n})\}$ of triples $(s_{i},t_{i},p_{i})$ indicating that $p_{i}$ fractional units of data are routed from $s_{i}$ to $t_{i}$, such that for all nodes $v$, $\sum_{i:s_{i}=v}p_{i}\leq 1$ and $\sum_{i:t_{i}=v}p_{i}\leq 1$. The relaxation to fractional matchings is at least defensible in settings where the motivation for restricting the transfers to be matchings arises from throughput constraints from the networking layer, as a fractional matching transfers the same amount of data per unit time as an integer matching. In contrast, this relaxation is probably not defensible in settings where the network is optical and the matching is being implemented by the rotation of mirrors [AS25a, Liu+23a, Mel+24a]. Thus we consider both allowing fractional matchings, and restricting matchings to be integral.
So the standard algorithmic research approach would be to, for each of these eight variations, find the best approximation ratio that is achievable by a polynomial-time algorithm. Thus let us start by noting what can be immediately deduced about the best achievable approximation ratio from known results. These observations are summarized in Table˜1.
fractional matching integral matching
direct indirect direct indirect
makespan 1 1 1 $O(\log n)$
average completion 1 $O(\log n)$ $\sqrt{2}$ [GM09a] $O(\log n)$
Table 1: Overview of known approximation guarantees.
For direct integral routing for the makespan objective, optimal schedules can be computed using an optimal algorithm for edge coloring a bipartite graph [Sch+03a]. For fractional routing with the makespan objective, the optimal makespan is the same for both direct and indirect routing because the fractional edge coloring number for a bipartite graph is the same as the maximum coloring number of the two underlying bipartite graphs [Sch+03a]. Finally it is straightforward to model fractional direct routing to optimize average completion time with a polynomial-size linear program. A $O(\log n)$ approximation is achievable for the rest of the variants by modeling the problem as a linear program, where the edges in each time step are over-capacitated by $O(\log n)$. It can be established that this over-capacitation is sufficient via the probabilistic method. For direct integral routing for average completion time, rounding the demands up to the next integer essentially does not change the problem, one can use the $\sqrt{2}$-approximation for integer demands is also $\sqrt{2}$-approximate for rational demands [GM09a].
In this paper, we focus on the variants where a $O(1)$ approximation does not (to the best of our knowledge) follow directly from known results, which all involve indirect routing.
1.2 Summary of Our Results
We present two main results. Theorem 1 presents the first $O(1)$-approximation algorithm for indirect fractional routing for the average completion time objective.
Theorem 1.
There is a polynomial-time indirect routing fractional matching algorithm $\mathrm{ALG}$} that outputs, for any valid instance, a schedule with an average completion time that is at most 16 times the optimal average completion time. The algorithm only uses direct routing schedules, so the approximation bound is valid both for the direct and indirect routing cases using fractional matchings.
The algorithm $\mathrm{ALG}$} is the natural greedy algorithm that routes an arbitrary maximal matching at each unit of time. Perhaps the limited utility of indirectness in the context of fractional routing for the objective of average completion time is somewhat surprising. Although recall that a similar situation arises for the makespan objective, where indirect routing is of no benefit.
Our algorithm introduces a novel dual fitting framework. As is usually the case in approximation algorithms, the critical step is to find the “right” lower bounds to the optimum to compare against. To that end, we introduce two relaxations of the problem. In the Sender Bound problem, there are only capacity constraints on the amount of data a sender can send in a unit of time, but there are no constraints on the amount of data a receiver can receive in a unit of time. Similarly, in the Receiver Bound problem, there are only capacity constraints on the amount of data a receiver can receive in a unit of time, but there are no constraints on the amount of data a sender can send in a unit of time. We consider the natural linear programming formulations PS and PR for the sender and receiver bound problems, and their respective duals DS and DR. We then use dual feasible solutions for DS and DR as our lower bounds. We are not aware of a similar “double” dual fitting analysis in the literature.
Our second main result appears in Theorem˜2, with a full characterization of the worst-case makespan and average completion time for indirect integral routing, with respect to two parameters $n$ and $B$. In the future, we plan to leverage the bounds in Theorem 2 towards determining whether a $O(1)$-approximation for indirect integral routing is possible.
Theorem 2.
Let $D$ be a rational demand matrix where the maximum row or column sum is at most $B$. There is a polynomial-time algorithm $\mathrm{ALG}$} that uses indirect integral routing, and guarantees makespan and average completion time
[…]
Further if each entry of $D$ is $\frac{B}{n}$ then the optimal schedule has makespan and average completion time $\Omega(\max\{\log n,\frac{B\log n}{\log B},B\})$. Thus, algorithm $\mathrm{ALG}$} is worst-case optimal with respect to both objectives.
Theorem 2 confirms the intuition that the worst-case input should be when the demand is evenly spread out. As $\left\lceil B\right\rceil$ is a clear lower bound on the optimal makespan (and optimal average completion time), we also derive the following corollary to the theorem:
Corollary 1.
Algorithm $\mathrm{ALG}$} from Theorem˜2 attains a $O(\log n)$-approximation ratio for both makespan and average completion time, for all $B$.
In some sense, the challenge of obtaining a $O(1)$-approximation appears to be to identify instances that admit schedules with less than worst-case objective value.
If $B\leq 2$, we can use Valiant’s hypercube design [Val82a] to achieve $O(\log n)$ makespan by using routing paths with $O(\log n)$ intermediate nodes. In a nutshell, this approach tries to maximize the usage of intermediate nodes, because there is very little demand to route. Accordingly, the $\Omega(\log n)$ makespan lower bound relies purely on topological properties, and does not take demands into account. Alternatively, if $B\geq n$, we can use a single round-robin design and route flow directly from sources to destinations in maximum matchings to achieve $O(B)$ makespan—there is no need for intermediate nodes because there are no small demands that can be “combined.” The $\Omega(B)$ lower bound relies purely on the maximum demand requested per source or destination node, without taking the network topology into account.
The challenge in the $2\leq B\leq n$ regime is thus to interpolate between these two extreme cases. For the upper bound, we use a generalization of the hypercube design, which uses $dn^{1/d}$ distinct matchings, for some dimension parameter $d$. Each matching must be repeated with a multiplicity $m$ that depends on $B,n,$ and $d$, and increases as $d$ increases, resulting in $mdn^{1/d}$ total makespan. Setting $d$ appropriately results in the desired makespan bound $O(B\log n/\log B)$. To prove the matching lower bound, we carefully partition routings into two categories. If a source can route to a large fraction of destinations using “short” paths with few intermediate nodes (even if only a small fraction of demand is routed there), we show a lower bound purely on network topology properties. Otherwise, if the routing uses “long” paths with many intermediate nodes for a large fraction of the total demand, we can prove a large makespan via a topology-independent averaging argument. This results in the desired lower bound $\Omega(B\log n/\log B)$ for an appropriate threshold of “short” versus “long” paths.
1.3 Related Work
Coflow scheduling was popularized by [CS12a], and followed by heuristics for online and offline settings [CZS14a, CS15a]. A recent overview of reconfigurable networking can be found in [AS25a].
There is some literature on direct integral routings to optimize weighted completion time. Marx [Mar09a] showed that this problem is APX-hard. The natural greedy algorithm is a 2-approximation [Bar+98a, Bar+00a]. [GM09a] give an improved $\sqrt{2}$-approximation, and [HKS11a] show a $1.83$-approximation for a more general setting. Other works in this area consider non-preemptive variants, where large demands must be scheduled consecutively in time [HKS01a, Kim05a, Gan+08a]. The fractional preemptive variants of these problems are also captured by the general polytope scheduling problem, as any feasible instantaneous allocation belongs to the matching polytope [IKM18a, JLM25a].
There is also some literature on the situation where there are several different applications, each with their own demand matrix, and the objective is to minimize the sum of the makespans over the different applications. This literature considers direct integral routing for the average completion time objective. The problem is APX-hard [Mar09a], and an even stronger inapproximability of better-than-2 follows from a reduction to the concurrent open shop problem [QS02a], which is hard under the Unique Games Conjecture [BK09a]. The first constant-factor approximation was shown by [QSZ15a], and after several improvements [KP16a, SG17a, Ahm+20a], the current best approximation ratio is $3.415$ [RS25a]. The special case where each coflow corresponds to the set of incident edges of a vertex is called data migration and also has been studied extensively [Kim05a, KKW04a, Gan+08a, GM09a, Mes10a]. Other notable results are online algorithms [Khu+19a, BNV20a, DM20a], a matroid generalization [Im+19a], and generalization for general graphs [Cho+19a, JKR17a]. In particular some elements of our dual fitting analysis derive from the dual fitting analysis in [DM20a].
The second main area of related work is in Oblivious Reconfigurable Networks (ORNs). The theoretical problem was formalized recently [Ami+22a], modeling the constraints of various proposed data center architectures using fast circuit switch technology [Gho+17a, Mel+24a, Shr+19a, AAS23a, Ami+24a]. The ORN model allows for multi-hop indirect routing, its key connection to our work. However, it also focuses on the streaming setting, where traffic demands arrive over time and must be continuously fulfilled. Additionally, they focus on obliviousness, and frame their results as a guarantee on the minimum amount of traffic which can be fulfilled, rather than as a guaranteed approximation ratio. One contribution of the ORN area is a generalized hypercube-style design with a tunable hop parameter [Ami+22a, Wil+23a, Ami+24a]. Our work in Section˜4 uses similar ideas, but applied to a different use case. Randomized oblivious connection schedules with some adaptive routing have also been considered in theory [Wil+24a], focusing on the same tradeoff guarantees as in the oblivous literature. It is still unknown what can be done with a fully adaptive design.
2 Notation and Formal Definitions
We follow broadly the definitions and notations used in [Ami+22a]. Our setting is a network with a collection $N=\{0,\ldots,n-1\}$ of $n$ nodes. The input is an $n\times n$ demand matrix $D$ with nonnegative rational entries, where $D_{ij}$ specifies the amount of data that needs to be transferred from node $i$ to node $j$.
To define feasible solutions for the problems, we consider a particular network flow instance derived from the demand matrix $D$. The network/graph $G$ has vertex set $N\times\{0,\ldots,T\}$. The edge set of $G$ is partitioned in edges $E_{\mathrm{phys}}$ that we will label as physical, and edges $E_{\mathrm{virt}}$ that we will label as virtual. The virtual edges are of the form $(i,t)\to(i,t+1)$ for all $i\in N$ and $t\in\{0,\ldots,T-1\}$. The physical edges are of the form $(i,t)\to(j,t+1)$ for some $i\neq j\in N$. Furthermore, we require the set of physical edges $M_{t}$ at time $t$ to form a matching in the complete bipartite graph given by node set $(\{(i,t)\}_{i\in N},\{(i,t+1)\}_{i\in N})$, i.e., each node $(i,t)$ and each node $(i,t+1)$ is the end point of at most one edge in $M_{t}$, for all $i\in N$ and for all $t\in\{0,\ldots,T-1\}$. For all $i,j\in N$ and $t\in\{1,\ldots,T\}$, let $\mathcal{P}(i,j,t)$ be the set of all directed paths from vertex $(i,0)$ to vertex $(j,t)$ in $G$. We define $\mathcal{P}:=\cup_{i,j\in N,t\geq 1}\mathcal{P}(i,j,t)$. An indirect fractinoal routing is a flow $f:\mathcal{P}\to\mathbb{R}_{\geq 0}$. We say that the routing (or flow) $f$ is feasible for demands $D$ if it satisfies the following constraints:
Capacity Constraints
For each edge $e\in E_{\mathrm{phys}}$, $\sum_{P\in\mathcal{P}}f(P)\cdot\mathbf{1}[e\in P]\leq 1$.
Demand Satisfaction
For each $i,j\in N$, $\sum_{t=1}^{T}\sum_{P\in\mathcal{P}(i,j,t)}f(P)\geq D_{ij}$.
If a routing $f$ routes $\alpha$ units of flow along a path $P$ from $(i,0)$ to $(j,t)$ that contains a physical edge $(a,s)\rightarrow(b,s+1)$ then $\alpha$ units of data with source $i$ and destination $j$ are transmitted from node $a$ to node $b$ at time $s$.
A routing $f$ is a direct routing if it sends positive flow only over paths that contain at most one physical edge. A routing $f$ is integral if every node has at most one incoming physical edge with positive flow, and at most one outgoing physical edge with positive flow at any time step. Note that with our definitions, direct and integral routings are a subset of indirect and fractional routings. A direct routing can be viewed as a sequence of fractional or integer matchings $M_{0},\ldots,M_{T-1}$ in the complete bipartite graph between two copies of the node set $N$ that covers $D$.
We say that a routing $f$ for the network $G$ has a makespan of $T$. The completion time $t$ for a unit of data with source $i$ and destination $j$ is the smallest $t$ such that this data arrives at a node of the form $(j,t)$. Thus, the total completion time of a schedule is then
$\sum_{t=1}^{T}\sum_{i,j\in N}\sum_{P\in\mathcal{P}(i,j,t)}t\cdot f(P)\ .$
3 Fractional Matching, Indirect Routing, and Average Completion Time
In this section we prove Theorem˜1. For clarity of explanation, we prove the theorem for the total sum of completion times, which is equivalent to the average (up to a factor of $1/\sum_{i,j}D_{ij}$). We start with a description of the polynomial-time algorithm $\mathrm{ALG}$} that gives us Theorem˜1.
Algorithm $\bm{{\mathrm{ALG}}}$ Description:
At each time $t$, the fractional matching $M_{t}$ is an arbitrary maximal fractional matching of the residual demand matrix $D(t)$.
The most important step in obtaining approximation ratio results is to find “good” lower bounds to the optimal objective value, which we will denote by $\mathrm{OPT}$}. Here we consider the following two lower bounds, each of which are derived by relaxing the constraints in two different (but symmetric) ways.
Sender Bound Problem Statement:
For each fractional matching $M=\{(s_{i},r_{i},p_{i})\}$ and for all nodes $v$, it must be the case that $\sum_{i:s_{i}=v}p_{i}\leq\frac{1}{4}$. However, $\sum_{i:r_{i}=v}p_{i}$ can be unbounded.
Receiver Bound Problem Statement:
For each fractional matching $M=\{(s_{i},r_{i},p_{i})\}$ and for all nodes $v$ it must be the case that $\sum_{i:r_{i}=v}p_{i}\leq\frac{1}{4}$. However, $\sum_{i:s_{i}=v}p_{i}$ can be unbounded.
In the sender bound problem there is no bound on the rate that data can be transferred into a node, and in the receiver bound problem there is no bound on the rate that data can be transferred out of a node. Instead of limiting the amount of flow that can be transferred into or out of a node to 1, as it may seem natural, we limit it to $\frac{1}{4}$. This setting arises due to our dual fitting argument. Note that
${\mathrm{OPT}}^{S}\leq 4\cdot{\mathrm{OPT}}\text{ and }{\mathrm{OPT}}^{R}\leq 4\cdot{\mathrm{OPT}}$ (1)
where ${\mathrm{OPT}}^{S}$ and ${\mathrm{OPT}}^{R}$ are optimal solutions to the sender bound and receiver bound problems, respectively. This is because replicating each matching in $\mathrm{OPT}$} four times results in a solution that is feasible for both the sender bound problem and the receiver bound problem, and increases the objective value by at most a factor of 4.
[…]
  • For all $i,j\in N$, we set $\alpha^{S}_{ij}=D_{i}^{S}(0)$ and $\alpha^{R}_{ij}=D_{j}^{R}(0)$.
  • For all $i,j\in N$, we set $\beta^{S}_{it}=\frac{1}{4}D^{S}_{i}(t)$ and $\beta^{R}_{jt}=\frac{1}{4}D^{R}_{j}(t)$.
[…]
Lemma 1.
$\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})\geq\frac{1}{2}{\mathrm{ALG}}$.
Proof.
Recall that
$\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})=\sum_{i,j\in N}D_{ij}\alpha^{S}_{ij}-\sum_{i,t}\beta^{S}_{it}+\sum_{i,j\in N}D_{ij}\alpha^{R}_{ij}-\sum_{j,t}\beta^{R}_{jt}\ .$
We first analyze $\mathrm{DS}(\alpha^{S},\beta^{S})$. Consider a particular sender receiver pair $(i,j)$. Let $T_{ij}$ denote the collection of times $t$ when $\mathrm{ALG}$} sends a positive amount of data from $i$ to $j$. If $\mathrm{ALG}$} does not send data from $i$ to $j$ at time $t$ it must be the case that either $\mathrm{ALG}$} is sending unit of data out of sender $i$ or into receiver $j$. Since there are at most $D_{i}^{S}+D_{j}^{R}-2D_{ij}$ such times, and sending all data from $i$ to $j$ takes at most $D_{ij}$ (fractional) time, for every $t\in T_{ij}$ it is the case that
$t\leq D_{i}^{S}+D_{j}^{R}-D_{ij}\ .$
Thus the completion time of any unit of data with source $i$ and destination $j$ is at most $D_{i}^{S}+D_{j}^{R}-D_{ij}$. This gives
${\mathrm{ALG}}$ $\leq\sum_{i,j\in N}D_{ij}\cdot(D_{i}^{S}+D_{j}^{R}-D_{ij})\leq\sum_{i,j\in N}D_{ij}\cdot(D_{i}^{S}+D_{j}^{R})=\sum_{i,j\in N}\left(D_{ij}\alpha_{ij}^{S}+D_{ij}\alpha_{ij}^{R}\right)\ .$
The equality follows by definition of the $\alpha$ variables. Now note that
$\frac{1}{2}{\mathrm{ALG}}$ $=\sum_{i\in N}D^{S}_{i}(t)+\sum_{j\in N}D^{R}_{j}(t)=\sum_{t=0}^{T}\sum_{i\in N}\beta^{S}_{it}+\sum_{t=0}^{T}\sum_{j\in N}\beta^{R}_{jt}$
The first equality follows because $\sum_{i}D^{S}_{i}(t)+\sum_{j}D^{R}_{j}(t)$ is equal to twice the total data at time $t$ that $\mathrm{ALG}$} has not yet transmitted. The second equality follows from the definition of $\beta$ variables. Combining these two bounds on $\mathrm{ALG}$}, we conclude
$\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})$ $=\sum_{i,j}D_{ij}\alpha^{S}_{ij}-\sum_{i,t}\beta^{S}_{it}+\sum_{i,j}D_{ij}\alpha^{R}_{ij}-\sum_{j,t}\beta^{R}_{jt}$
$\geq{\mathrm{ALG}}-\frac{1}{2}{\mathrm{ALG}}=\frac{1}{2}{\mathrm{ALG}}\ .$
This completes the proof of the lemma. ∎
Lemma 2.
[…]
Proof.
[…]
We now conclude the proof of Theorem˜1:
Proof of Theorem˜1.
We now conclude by noting
$4\cdot{\mathrm{OPT}}+4\cdot{\mathrm{OPT}}$ $\geq{\mathrm{OPT}}^{R}+{\mathrm{OPT}}^{S}\geq\mathrm{DS}(\alpha^{S},\beta^{S})+\mathrm{DR}(\alpha^{R},\beta^{R})\geq\frac{1}{2}\cdot{\mathrm{ALG}}\ .$
The first inequality follows from the observation in Line (1). The second inequality follows from weak duality and the dual feasibility (from Lemma˜2). Finally, the last inequality follows from Lemma˜1. Thus we can conclude that the approximation ratio of $\mathrm{ALG}$} is at most $16$. ∎
4 Integral Matching and Indirect Routing
In this section, we present the proof of Theorem˜2. We first consider the special case where every entry of the demand matrix is $B/n$, and show worst-case upper and lower bounds on makespan and average completion time. At the end of this section, we will then show that we can lift these results to the general setting while only losing small constants.
We distinguish three main cases, $B>n$, $B<2$, and $2\leq B\leq n$, which we handle in the following. All of the proposed algorithms below run in polynomial time.
Lemma 3.
If all demands are equal to $B/n$ and $B>n$, then there exists an indirect integral routing scheme with makespan and average completion time at most $O(B)$. Moreover, every indirect integral routing scheme has makespan and average completion time at least $\Omega(B)$.
Proof.
We use a single round-robin design, which is in fact a direct routing scheme. However, we will see that it achieves a best-possible makespan in the worst-case (up to a constant) even among indirect routing schemes.
The connection schedule consists of $n$ distinct matchings, equal to the set of all cyclic permutations of $n$, with each matching repeated $\left\lceil\frac{B}{n}\right\rceil$ times. With this connection schedule, every possible source-destination pair is directly connected with multiplicity $\left\lceil\frac{B}{n}\right\rceil$. The routing protocol chooses to always route flow on direct connections. This gives us total makespan equal to the total number of matchings, $\left\lceil\frac{B}{n}\right\rceil n\leq 2B$.
To prove the lower bound, note that every indirect routing scheme must be able to route all demands out of every source, and into every destination. In particular, the maximum row or column sum of the demand matrix is a lower bound on the makespan, and the maximum row or column sum is exactly $B$ when all demands are equal to $B/n$. Moreover, the median completion time of demands transmitted by any sender is at least $B/2$, and therefore, the average completion time of all demands is $\Omega(B)$. ∎
Lemma 4.
If all demands are equal to $B/n$ and $B<2$, then there exists an indirect integral routing scheme with makespan and average completion time at most $O(\log n)$. Moreover, every indirect integral routing scheme has makespan and average completion time at least $\Omega(\log n)$.
Proof.
We use a variant of Valiant’s hypercube design [VB81a]. Assume w.l.o.g. that $n$ is a power of $2$, and represent each node $u\in\{1,\ldots,n\}$ in binary. That is, $u=(u_{1},\ldots,u_{\log_{2}(n)})$ where $u_{i}\in\{0,1\}$ for all $i$. The connection schedule $\mathcal{M}$ will consist of exactly $\log_{2}(n)$ matchings. For each $i\in\{1,\ldots,\log_{2}(n)\}$, the $i$th matching connects nodes which match in all but the $i$th coordinate. That is,
$M_{i}(u_{1},\ldots,u_{\log_{2}(n)})=(u_{1},\ldots,u_{i-1},u_{i}\oplus 1,u_{i+1},\ldots,u_{\log_{2}(n)}).$
The routing protocol routes flow on the shortest path from source to destination. That is, to route from $u$ to $v$, a single physical edge is taken for every index for which $u$ and $v$ do not match. This strategy obtains makespan $\log_{2}(n)$. All that is left is to show that this routing protocol does not overload the physical edges from our matchings.
Consider a node’s binary representation. On average each node $u$ will agree with other nodes $v$ on half of their binary indices. Therefore every node pair $u,v$ is connected by a routing path with an average of $\log_{2}(n)/2$ physical edges. Because demands are uniform, all flow in the network travels along $\log_{2}(n)/2$ physical edges on average. Additionally, this also implies that the total flow $Bn$ will be evenly load-balanced across all $n\log_{2}(n)$ physical edges of the network, leading to exactly
$\frac{\log_{2}(n)}{2}\cdot\frac{Bn}{n\log_{2}(n)}\leq 1$
flow traversing each physical edge, using $B\leq 2$. Thus, the flow is feasible on $\mathcal{M}$.
To prove the lower bound, note that in order to route uniform demands $B/n$ for any $B>0$, there must be some routing path between every source-destination pair $u,v$. This is only possible when the number of distinct paths which leave any fixed node $u$ is at least $n$. Given some makespan bound $L$, the number of distinct paths leaving a fixed node $u$ using exactly $1\leq d\leq L$ physical edges is equal to $L\choose d$}. Thus, the number of distinct paths leaving $u$ using at most $L$ physical edges is exactly $\sum_{d=1}^{L}{L\choose d}=2^{L}-1$. Therefore, it must be the case that $n\leq 2^{L}$, or equivalently, $L\geq\log_{2}(n)$. Using the same argument on source $u$ for reaching $n/2$ destinations, we can see the median completion time of demands send from $u$ is at least $\log_{2}(n)/2$, hence the average completion time of all demands is $\Omega(\log n)$. ∎
For the case $2\leq B\leq n$, we split our results into two lemmas. We start with the upper bounds.
Figure 2: The time-expanded routing scheme of a hypercube with dimension $d=2$ and $n=9$.
Lemma 5.
If all demands are equal to $B/n$ and $2\leq B\leq n$, then there exists an indirect integral routing scheme with makespan and average completion time at most $O(B\log n/\log B)$.
Proof.
Let $d=\left\lceil\log_{2}n/\log_{2}B\right\rceil$ and assume w.l.o.g. that $n^{1/d}$ is an integer. Represent nodes $u\in\{0,\ldots,n-1\}$ as $d$-tuples $u=(u_{1},\ldots,u_{d})$, where each $u_{i}\in\{0,\ldots,n^{1/d}-1\}$. For our connection schedule, we use a single period of the Elementary Basis connection schedule (Section 4.1 of [Ami+22a]), illustrated in Figure˜2 for dimension parameter $d=2$ and explained in detail below. We will use multiplicity of matchings $\left\lceil B/n^{1/d}\right\rceil$. For each index $i\in\{1,\ldots,d\}$ and each scale factor $s\in\{1,\ldots,n^{1/d}-1\}$, we define the matching $M_{t=(i,s)}$ which connects nodes $u$ to nodes $v$ which match $u$ in all but the $i$th coordinates, and which differ from $u$ in the $i$th coordinate by exactly $s(\mbox{mod }n^{1/d})$. That is,
$M_{t=(i,s)}(u)=(u_{1},\ldots,u_{i-1},u_{i}+s(\mbox{mod }n^{1/d}),u_{i+1},\ldots,u_{d}).$
Our connection schedule repeats each matching $M_{t}$ exactly $\left\lceil B/n^{1/d}\right\rceil$ times. Hence, the length of the schedule is equal to $T=d\cdot(n^{1/d}-1)\cdot\left\lceil B/n^{1/d}\right\rceil$.
To route between an arbitrary node pair $u$ to $v$, greedily route flow across any physical edge which connects to a node with decreased Hamming distance to $v$ (i.e. it matches $v$ in the modified $i$th coordinate). This process will use at most $d$ physical edges to route between each pair $u,v$.
We next show that no edge is overloaded. Since there are $n^{1/d}$ choices for each of the $d$ entries of node’s tuples, on average each node $u$ agrees with an other node $v$ on $d/n^{1/d}$ entries. Therefore, every node pair $u,v$ is connected by a routing path using an average of $d-d/n^{1/d}$ physical edges. Because uniform traffic is being routed, flow will be evenly balanced across all physical edges. Since a total of $Bn$ flow is routed over $Tn$ physical edges, the load on each edge is
$\bigg(d-\frac{d}{n^{1/d}}\bigg)\cdot\frac{Bn}{Tn}=\bigg(d-\frac{d}{n^{1/d}}\bigg)\cdot\frac{B}{d\cdot(n^{1/d}-1)\cdot\left\lceil\frac{B}{n^{1/d}}\right\rceil}=\frac{B}{n^{1/d}\cdot\left\lceil\frac{B}{n^{1/d}}\right\rceil}\leq 1\ .$
Finally, we can bound the makespan as follows:
$T$ $=\left\lceil\frac{B}{n^{1/d}}\right\rceil\cdot dn^{1/d}\leq dn^{1/d}+dB=\left\lceil\frac{\log_{2}n}{\log_{2}B}\right\rceil\cdot\left(n^{1/\left\lceil\frac{\log_{2}n}{\log_{2}B}\right\rceil}+B\right)$
$\leq\left(\frac{\log_{2}n}{\log_{2}B}+1\right)\cdot\left(n^{\frac{\log_{2}(B)}{\log_{2}(n)}}+B\right)=\left(\frac{\log_{2}n}{\log_{2}B}+1\right)\cdot 2B\in\mathcal{O}\left(\frac{B\log_{2}(n)}{\log_{2}(B)}\right).$
This completes the proof of the lemma. ∎
Lemma 6.
If all demands are equal to $B/n$ and $2\leq B\leq n$, then every indirect integral routing scheme has makespan and average completion time at least $\Omega(B\log n/\log B)$.
Proof.
Fix any integral routing scheme $f$ and let $T$ be its makespan. For any node pair $u,v$, let $S_{u,v}$ denote the minimum number of physical edges used to route demands from $u$ to $v$, and let $h$ be the median of $S_{u,v}$ over all node pairs $u,v$. Let $d=\frac{1}{3}\log_{2}n/\log_{2}B$. We distinguish whether $h\leq d$ or $h\leq d$.
Case 1: $h\geq d$. Because we assumed $f$ obtained makespan $T$, there are exactly $nT$ total physical edges that may be used to route flow. Additionally, there are $Bn$ total units of flow that need to be routed across the entire network. In particular, each node pair $u,v$ with $S_{u,v}\geq h$ requires at least $h$ physical edges per unit flow. Thus, there is a total of $\frac{1}{2}Bn$ demand that needs at least $h$ physical edges per unit flow. Thus, the average amount of flow per physical edge is at least $\frac{Bnh}{2nT}\geq\frac{Bd}{2T}$ . Since each physical edge has unit capacity and we assumed that $f$ was feasible, it must be that $T\geq Bd/2\in\Omega(B\log_{2}(n)/\log_{2}(B))$.
Case 2: $h\leq d$. By the definition of $h$, for half of the node pairs $f$ routes some flow using at most $h$ physical edges. Thus, there must exists a sender $u^{*}$ that can route flow to at least $n/2$ receivers using at most $h$ physical edges each. Therefore, there must exist at least $n/2$ distinct paths from $u^{*}$ that reach those receivers within makespan $T$ using at most $h$ physical edges. We next compute how many such paths exist. There are $T\choose i$} distinct paths using exactly $i$ physical edges that obtain makespan $T$. Thus, there are $\sum_{i=1}^{h}{T\choose i}$ distinct paths with at most $h$ physical edges within makespan $T$. Thus, it must be that $n/2\leq\sum_{i=1}^{h}{T\choose i}$.
In the following, we derive a lower bound on $T$ from this inequality. We first note that the proof of Lemma˜4 shows that for any $B>0$, it holds that $T\geq\log_{2}n$, and thus, $T\geq\log_{2}n/\log_{2}B=3d$ since we assume that $B\geq 2$. We can now observe that
$\sum_{i=1}^{h}{T\choose i}\leq 2{T\choose h}\leq 2\frac{T^{h}}{h!}\ ,$
where the first inequality holds using $h\leq d\leq T/3$, and the second inequality is a well-known approximation of the binomial coefficient. Rearranging $n/2\leq 2T^{h}/h!$ gives us
$T\geq\bigg(h!\cdot\frac{n}{4}\bigg)^{1/h}=n^{1/h}\cdot(h!)^{1/h}\cdot\bigg(\frac{1}{4}\bigg)^{1/h}\ .$
Using Stirling’s approximation $h!\geq\sqrt{2\pi h}\cdot(\frac{h}{e})^{h}$, we can derive $T\geq\frac{h}{e}n^{1/h}\cdot\left(\frac{\sqrt{2\pi h}}{4}\right)^{1/h}\geq\frac{h}{e}n^{1/h}\ ,$ because $\left(\frac{\sqrt{2\pi h}}{4}\right)^{1/h}$ converges to one from above as $h\to\infty$.
The derivative of $h\cdot n^{1/h}$ is equal to $n^{1/h}(h-\ln n)/h$, which is negative for all $2\leq h<\ln n$. Hence, the function $h\cdot n^{1/h}$ is strictly decreasing on this interval. Since $\ln n>\frac{1}{2}\log_{2}n\geq\frac{1}{2}\log_{2}n/\log_{2}B\geq d$, we can conclude that $h\cdot n^{1/h}\geq d\cdot n^{1/d}$ for all $2\leq h\leq d$. Thus, $f$ has makespan at least $T\geq h\cdot n^{1/h}/e\in\Omega(B\log_{2}(n)/\log_{2}(B))$. This completes the proof of Case 2.
Note that by our choice of $h$, our lower bounds on the makespan also apply to the median completion time, hence also give the same lower bound on the average completion time up to constants. This completes the proof of the lemma. ∎
This completes our results for the uniform demand setting. To lift our results to the general setting, we use the following well-known reduction.
Lemma 7.
Given a connection schedule $\mathcal{M}=M_{0},\ldots,M_{T-1}$ and routing protocol $f_{unif}$ over $\mathcal{M}$ that is feasible for uniform demands $B/n$ and obtains makespan $T$, then for any demand matrix $D$ with row and column sums no more than $B$, we can build a routing $f_{D}$ over the connection schedule $\mathcal{M}\circ\mathcal{M}$ that obtains makespan $2T$.
Proof.
This is a natural application of Valiant Load Balancing [VB81a], which is described at a high level below. To route $r$ demand between any node pair $u,v$, first traffic is routed from $u$ to all intermediate nodes $i$ in the network in equal shares $r/n$, using $f_{unif}$. Then, traffic is routed from intermediate nodes to its destination using $f_{unif}$. This allows us to treat routing any demand $D$ with row and column sums no more than $B$ as if it is two copies of uniform demand, requesting $B/n$ demand between node pairs. ∎
5 Conclusion
In this paper, we focus on two of the variants of coflow scheduling with rational demand matrices, for which $O(1)$-approximations were not previously known: Namely, we present a 16-approximation algorithm for minimizing the average completion time under indirect fractional matching routings. Secondly, we address both the makespan and the average completion time objectives under indirect integral routing, and provide a full characterization of the worst-case runtimes under both objectives for this regime. While this falls short of providing $O(1)$-approximations for indirect integral routing, our proposed algorithm provides $O(\log n)$-approximations for the two objectives, in addition to being worst-case optimal; we also believe that the worst-case analysis will be helpful in determining whether a $O(1)$-approximation for indirect integral routing is possible. If it is, then we will have shown, together with the results in this paper, that constant-approximation algorithms exist for all possible coflow scheduling variants in Table˜1.
{refcontext}
[sorting=nyt]